OpenStudy (anonymous):

Use the general power rule to evaluate the integral:

OpenStudy (anonymous):

\[\int\limits_{}^{}x \sqrt{8-4x ^{2}} dx\]

OpenStudy (shubhamsrg):

if you see properly, if you let whole of what is inside the sq. root , as t , then a very easy expression to integrate shall come into play..

OpenStudy (anonymous):

i know just the general way of solving it in which u have to substitute \[4x ^{2}=t\] differentiate both sides wrt to x 8x=dt/dx (8x)dx=dt the integral after substitution changes to \[(1/8)\int\limits_{}^{}\sqrt{8-t}dt\]

OpenStudy (anonymous):

ok! remember to close the question guys and if appropriate, to award medals:)